### v1

@@ Abstract
We illustrate the approach by reanalyzing the Baxter colorectal adenoma dataset and comparing our results with Greenacre's ordination-based analysis using Correspondence Analysis and Canonical Correspondence Analysis. Our models consistently recover a core subset of taxa previously identified as discriminant, thereby corroborating Greenacre's main findings, while also revealing additional OTUs that become important once demographic covariates are taken into account. In particular, adjustment for age, gender, and diabetes medication improves the precision of the separation index and highlights new, potentially relevant taxa, suggesting that part of the original signal may have been influenced by confounding.
@@
Compositional data analysis, Microbiome, Multinomial classification, OTU separability.
@@ 1 Introduction
High-throughput sequencing technologies have made it possible to characterize microbial communities in great detail, typically through counts of operational taxonomic units (OTUs) [4]. The resulting data are inherently compositional: each sample is constrained to a fixed total, a property known as the “closure” constraint ([1], [9]). As a consequence, microbiome datasets generated by high-throughput sequencing should be treated as compositional at every stage of the analysis ([9], [15]). When standard multivariate methods are applied directly to raw or proportion-normalized counts, they can introduce spurious associations and lead to misleading inferences [9]. This issue is particularly problematic in supervised learning, where the goal is often to identify a subset of OTUs that best discriminates between multiple phenotypic or clinical groups.
@@
The challenge of identifying discriminative features from compositional data is not unique to microbiome research. Compositional structures arise in many other fields, including geochemistry, metabolomics, ecology, single-cell transcriptomics, and economics, whenever measurements represent parts of a whole [8]. In all these settings, statistical and machine learning methods that ignore the compositional constraint risk producing biased estimates and incorrect interpretations [17]. There is therefore a clear need for classification and feature-selection frameworks that explicitly account for compositionality, in order to support valid inference, interpretability, and reproducibility in high-dimensional applications.
@@
In microbiome studies in particular, researchers are often interested not only in accurate prediction but also in interpretable variable selection identifying “separable” OTUs whose relative abundances differ systematically across groups and that can be combined into discriminative microbial signatures. Existing approaches typically rely on marginal testing, methods developed for binary outcomes, or procedures that do not fully respect the compositional structure of the data [18]. Moreover, they rarely offer a coherent way to quantify uncertainty in the selected OTUs or to propagate OTU-level variability into global measures of separability in multinomial classification problems.
@@
In this work, we introduce a classification framework specifically designed for compositional microbiome data, with the aim of identifying microbial taxa that most effectively discriminate among multiple biological or clinical groups. The framework starts with standard preprocessing steps to ensure data quality and comparability, including the removal of rare taxa, Bayesian-multiplicative imputation of zeros, and centered log-ratio transformation to properly handle compositional constraints ([15]; |citeGloor2017).
@@
At its core, the methodology uses a regression-based model that relates microbial composition to group membership while accounting for the inherent interdependence among components. Extending this model to the multinomial setting allows us to work naturally with more than two groups. By expressing the model in terms of log-ratios between pairs of OTUs, we can directly examine how the balance between taxa contributes to group differentiation ([10], [12]).
@@
Finally, we incorporate formal procedures to quantify the uncertainty associated with the separability estimates, combining analytical variance estimators with resampling-based inference. This allows us to ensure that the OTUs we identify are not only discriminative but also statistically robust. The complete framework is implemented in the codabiocom R package, which offers tools for covariate adjustment, uncertainty quantification, and parallel computation, making it suitable for large-scale compositional datasets.
@@ 2 Preliminaries
Microbiome datasets obtained through high-throughput sequencing consist of abundance counts for $m$ operational taxonomic units (OTUs) across $n$ samples. These data are compositional by nature: only relative abundances are meaningful, and the total abundance within each sample is constrained to a fixed sum. Throughout this section, we describe the preprocessing steps and fundamental definitions used in our framework for OTU classification and selection.
@@ 2.1 Data Preprocessing
To ensure robustness and comparability across samples, the abundance matrix $\mathbf{X}$ of dimension $n\times m$ undergoes a series of preprocessing steps:
@@
Because compositional data analysis (CoDA) relies on logarithmic transformations, zero counts must be replaced with small positive values. We employ the Bayesian-multiplicative imputation proposed by Martin-Fernandez et al. [16], which preserves the sample composition while maintaining relative ratios among non-zero components.
@@
The first step of the workflow (Step 1 in Figure 1) involves removing low-frequency OTUs, imputing zeros, and applying the centered log-ratio (clr) transformation.
2.2 Basic Definitions
For clarity, we introduce terminology used throughout the paper.
Differential OTUs.
A subset $S=\{\text{OTU}_{i_{1}},\ldots,\text{OTU}_{i_{k}}\}$ is said to be differential if it provides measurable separation between two or more biological groups, according to a specified separability criterion.
Separability measure.
For a given subset $S$, we define a separability index $\mathrm{Sep}(S)$ as
$\mathrm{Sep}(S)=\frac{h}{\sqrt{h_{1}^{2}+h_{2}^{2}}},$
where $h$ denotes the height of the dendrogram at which two groups become separated and $h_{i}$ is the within-group height for group $i$. This index quantifies how distinctly the two clusters differ.
A subset $S_{1}$ is considered superior to $S_{2}$ if $\mathrm{Sep}(S_{1})>\mathrm{Sep}(S_{2})$.
@@ 3 Binary classification framework
We first describe our methodology in the binary setting, which provides the basis for its extension to multinomial classification. Let $\mathbf{X}$ be the $n\times m$ abundance matrix, with rows corresponding to samples and columns to OTUs, and let $Y$ denote the group indicator taking values in $\{0,1\}$. Our aim is to identify those OTUs (or combinations thereof) that best discriminate between the two groups, while respecting the compositional nature of the data.
@@ 3.1 CoDA-based regression for microbiome analysis
We address the classification problem using a penalized regression model in which discriminative OTUs are those with nonzero regression coefficients. More specifically, we build on the microbiome signature based on log-ratio analysis proposed in Calle et al. [5].
@@
subject to the zero-sum constraint $\sum_{j=1}^{m}\alpha_{j}=0$. This constraint enforces the scale-invariance principle required in compositional data analysis (CoDA), ensuring that the model depends only on log-ratios among components rather than on their absolute scale.
@@
Variable selection is then performed by fitting a penalized regression model to (1). The regression coefficients $\theta_{ij}$ are estimated by minimizing a loss function $L(\theta)$ with an elastic-net penalty:
@@
The penalty parameters $\lambda_{1}$ and $\lambda_{2}$ can be reparametrized as
@@
where $\mathbf{M}$ is the matrix of all pairwise log-ratios, of dimension $n\times\frac{m(m-1)}{2}$. Extensions to logistic regression and other generalized linear models are discussed in Friedman et al. [7] and apply directly in our setting.
@@ 3.2 Pairwise log-ratio screening and separability index
    • For multiclass classification ($C>2$), we employ the generalized AUC proposed by Hand and Till [13],
@@ 3.3 Uncertainty quantification
    • Binary case ($C=2$). Two common approaches exist for estimating $\operatorname{Var}(\widehat{AUC}_{jj^{\prime}})$.
      First, the classical approximation of Hanley and McNeil [11]:
      $\operatorname{Var}_{\text{Hanley}}(\widehat{AUC}_{jj^{\prime}})$ $=\frac{AUC_{jj^{\prime}}(1-AUC_{jj^{\prime}})+(n_{1}-1)(Q_{1}-AUC_{jj^{\prime}}^{2})}{n_{1}n_{0}}+$
      $\frac{(n_{0}-1)(Q_{2}-AUC_{jj^{\prime}}^{2})}{n_{1}n_{0}},$ (3)
      $Q_{1}$ $=\frac{AUC_{jj^{\prime}}}{2-AUC_{jj^{\prime}}},\quad Q_{2}=\frac{2\,AUC_{jj^{\prime}}^{2}}{1+AUC_{jj^{\prime}}}.$
      Second, the nonparametric U-statistic approach of DeLong et al. [6], for which a fast implementation is given by Sun and Xu [19]. Define
@@
      where $AUC_{cc^{\prime}}$ is the binary AUC comparing classes $c$ and $c^{\prime}$. The variance of $AUC_{jj^{\prime}}$ then propagates the uncertainty of each binary component:
@@
      Here, $\operatorname{Var}(AUC_{cc^{\prime}})$ can be approximated using the binary formula in (1), while the covariances between binary pairs need to be approximated or estimated.
  2. 2.
    Dependence among AUCs across OTU pairs. Each $AUC_{jj^{\prime}}$ shares OTUs with other pairs, inducing dependence in the matrix $\mathbf{A}$. When aggregating AUCs to compute the separability index $S_{k}$, this dependence propagates to its variance.
@@
    Pairs that share at least one OTU (e.g., $(AUC_{12},AUC_{13})$) are typically positively correlated. In practice, these covariances are non-negligible but difficult to derive analytically. A feasible approximation is
    $\operatorname{Cov}(AUC_{jj^{\prime}},AUC_{ll^{\prime}})\approx\rho_{OTU}\sqrt{\operatorname{Var}(AUC_{jj^{\prime}})\operatorname{Var}(AUC_{ll^{\prime}})}$
    if the pairs share at least one OTU, where $\rho_{OTU}$ can be estimated empirically from resampling runs or set to a conservative value (e.g., 0.1–0.3).
    For the multiclass case, an additional layer of dependence arises because the Hand–Till AUC is itself an average of dependent binary pairs. An analogous approximation can be used for $\operatorname{Cov}(AUC_{cc^{\prime}},AUC_{ll^{\prime}})$ when the class pairs share at least one class.
  3. 3.
    Alternative: non-parametric bootstrap. As an alternative to analytic variance formulas, we consider a non-parametric bootstrap procedure. This approach is particularly useful in capturing the sampling variability of $S_{k}$ under complex dependency structures:
@@
Objective: Identify differential OTUs that best discriminate between groups $(g=1,\ldots,k)$ using CoDA and pairwise log-ratio modeling. Input data: Abundance matrix $\mathbf{X}_{n\times m}$ where rows = samples, columns = OTUs, and class labels $Y\in\{1,\ldots,k\}$. 1. Data Preprocessing:

### v2

@@ Abstract
We illustrate the approach by reanalyzing the Baxter colorectal adenoma dataset and comparing our results with Greenacre’s ordination-based analysis using Correspondence Analysis and Canonical Correspondence Analysis. Our models consistently recover a core subset of taxa previously identified as discriminant, thereby corroborating Greenacre’s main findings, while also revealing additional OTUs that become important once demographic covariates are taken into account. In particular, adjustment for age, gender, and diabetes medication improves the precision of the separation index and highlights new, potentially relevant taxa, suggesting that part of the original signal may have been influenced by confounding.
@@
Compositional data analysis; Microbiome; Multinomial classification; OTU separability.
@@ 1 Introduction
High-throughput sequencing technologies have made it possible to characterize microbial communities in great detail, typically through counts of operational taxonomic units (OTUs) (Caporaso et al., 2011). The resulting data are inherently compositional: each sample is constrained to a fixed total, a property known as the “closure” constraint (Aitchison, 1986; Gloor et al., 2017). As a consequence, microbiome datasets generated by high-throughput sequencing should be treated as compositional at every stage of the analysis (Gloor et al., 2017; Fernandes et al., 2014). When standard multivariate methods are applied directly to raw or proportion-normalized counts, they can introduce spurious associations and lead to misleading inferences (Gloor et al., 2017). This issue is particularly problematic in supervised learning, where the goal is often to identify a subset of OTUs that best discriminates between multiple phenotypic or clinical groups.
@@
The challenge of identifying discriminative features from compositional data is not unique to microbiome research. Compositional structures arise in many other fields, including geochemistry, metabolomics, ecology, single-cell transcriptomics, and economics, whenever measurements represent parts of a whole (Greenacre, Martínez-Álvaro & Blasco, 2021). In all these settings, statistical and machine learning methods that ignore the compositional constraint risk producing biased estimates and incorrect interpretations (Quinn et al., 2018). There is therefore a clear need for classification and feature-selection frameworks that explicitly account for compositionality, in order to support valid inference, interpretability, and reproducibility in high-dimensional applications.
@@
In microbiome studies in particular, researchers are often interested not only in accurate prediction but also in interpretable variable selection—identifying “separable” OTUs whose relative abundances differ systematically across groups and that can be combined into discriminative microbial signatures (Susin et al., 2020). Existing approaches typically rely on marginal testing, methods developed for binary outcomes, or procedures that do not fully respect the compositional structure of the data (Susin et al., 2020). Moreover, they rarely offer a coherent way to quantify uncertainty in the selected OTUs or to propagate OTU-level variability into global measures of separability in multinomial classification problems.
@@
In this work, we introduce a classification framework specifically designed for compositional microbiome data, with the aim of identifying microbial taxa that most effectively discriminate among multiple biological or clinical groups. The framework starts with standard preprocessing steps to ensure data quality and comparability, including the removal of rare taxa, Bayesian-multiplicative imputation of zeros, and centered log-ratio transformation to properly handle compositional constraints (Martín-Fernández et al., 2015; Gloor et al., 2021).
@@
At its core, the methodology uses a regression-based model that relates microbial composition to group membership while accounting for the inherent interdependence between components. Extending this model to the multinomial setting allows us to work naturally with more than two groups. By expressing the model in terms of log-ratios between pairs of OTUs, we can directly examine how the balance between taxa contributes to group differentiation (Greenacre et al., 2021; Hinton et al., 2022).
@@
Finally, we incorporate formal procedures to quantify the uncertainty associated with separability estimates, combining analytical variance estimators with resampling-based inference. This allows us to ensure that the OTUs we identify are not only discriminative but also statistically robust. The complete framework is implemented in the codabiocom R package [4], which offers tools for covariate adjustment, uncertainty quantification, and parallel computation, which makes it suitable for large-scale compositional datasets.
@@ 2 Preliminaries
The microbiome datasets obtained through high-throughput sequencing consist of abundance counts for operational taxonomic units (OTUs) $m$ in $n$ samples. These data are compositional by nature: only relative abundances are meaningful, and the total abundance within each sample is constrained to a fixed sum. Throughout this section, we describe the preprocessing steps and fundamental definitions used in our framework for OTU classification and selection.
@@ 2.1 Data Preprocessing
To ensure robustness and comparability between samples, the abundance matrix $\mathbf{X}$ of dimension $n\times m$ undergoes a series of preprocessing steps:
@@
Because compositional data analysis (CoDA) relies on logarithmic transformations, zero counts must be replaced with small positive values. We employ the Bayesian-multiplicative imputation proposed by [11], which preserves the sample composition while maintaining relative ratios among non-zero components.
@@
The first step of the workflow (Step 1 in Figure 1) involves removing low-frequency OTUs, imputing zeros and applying the centered log-ratio (clr) transformation.
2.2 Concept of separable OTUs
In this paper we use the term separable OTUs to denote taxa that, taken together, allow for a good discrimination between the outcome classes in a compositional setting. Let $\mathcal{G}=\{1,\ldots,G\}$ denote the set of classes and let $\mathcal{J}=\{1,\ldots,m\}$ index the OTUs. For any pair of OTUs $j,j^{\prime}\in\mathcal{J}$, we consider the log-ratio
$z_{ijj^{\prime}}=\log\frac{x_{ij}}{x_{ij^{\prime}}},$
where $\mathbf{x}_{i}=(x_{i1},\ldots,x_{im})$ denotes the composition of sample $i$. A pair $(j,j^{\prime})$ is said to be *separable* for a given classification task if the distribution of $z_{ijj^{\prime}}$ differs substantially across outcome classes.
To quantify separability we fit, for each ordered pair $(j,j^{\prime})$, a (multinomial) logistic regression model with $z_{ijj^{\prime}}$ as predictor and the outcome as response, and compute the corresponding generalized AUC. This yields a matrix of pairwise AUC values, which we denote by
$A=\bigl(\text{AUC}_{jj^{\prime}}\bigr)_{j,j^{\prime}=1}^{m}.$
High values of $\text{AUC}_{jj^{\prime}}$ indicate that the log-ratio between OTUs $j$ and $j^{\prime}$ provides good discrimination between classes; low values indicate limited discriminatory power.
3 Separability index for a set of OTUs
Our goal is to rank OTUs not only individually, but also in terms of how well *sets* of OTUs perform in combination. For any subset $S\subseteq\mathcal{J}$ with $|S|=k$, we define its *separability index* as
$\mathrm{Sep}(S)=\frac{1}{k(k-1)}\sum_{j\in S}\sum_{\begin{subarray}{c}j^{\prime}\in S\\ j^{\prime}\neq j\end{subarray}}\text{AUC}_{jj^{\prime}},$
that is, the average of all pairwise AUC values associated with the OTUs in $S$. Intuitively, a set $S$ will have a high separability index if most of its log-ratio pairs induce good discrimination between the outcome classes.
In practice, we construct a ranking of OTUs, $j_{(1)},j_{(2)},\ldots,j_{(m)}$, where $j_{(1)}$ denotes the most individually discriminant OTU, $j_{(2)}$ the second most discriminant, and so on (see Section X.X for details). For each $k=2,\ldots,k_{\max}$ we then consider the top-$k$ set
$S_{k}=\{j_{(1)},\ldots,j_{(k)}\}$
and compute its separability index
$S_{k}=\mathrm{Sep}(S_{k}).$
The curve $k\mapsto S_{k}$ describes how average pairwise separability improves (or deteriorates) as more OTUs are included. We choose the number of OTUs $\hat{k}$ as the value maximizing $S_{k}$,
$\hat{k}=\arg\max_{k}S_{k},$
and refer to the OTUs in $S_{\hat{k}}$ as the *separable OTUs* for the classification task under study.
This definition is fully determined by the AUC matrix $A$ and does not rely on any particular clustering or dendrogram representation. We therefore focus on $S_{k}$ as our main separability index throughout the rest of the paper.
@@ 4 Binary classification framework
We first describe our methodology in the binary setting, which forms the basis for its extension to multinomial classification. Let $\mathbf{X}$ be the $n\times m$ abundance matrix, with rows corresponding to samples and columns corresponding to OTUs, and let $Y$ denote the group indicator taking values in $\{0,1\}$. Our goal is to identify those OTUs (or combinations thereof) that best discriminate between the two groups, while respecting the compositional nature of the data.
@@ 4.1 CoDA-based regression for microbiome analysis
We address the classification problem using a penalized regression model in which discriminative OTUs are those with nonzero regression coefficients. More specifically, we build on the microbiome signature based on log-ratio analysis proposed in [4].
@@
subject to the zero-sum constraint $\sum_{j=1}^{m}\alpha_{j}=0$.
This constraint enforces the scale-invariance principle required in compositional data analysis (CoDA), ensuring that the model depends only on log-ratios among components rather than on their absolute scale.
@@
Variable selection is then carried out by fitting a penalized regression model to (1). The regression coefficients $\theta_{ij}$ are estimated by minimizing a loss function $L(\theta)$ with an elastic-net penalty:
@@
The penalty parameters $\lambda_{1}$ and $\lambda_{2}$ can be reparameterized in terms of $(\lambda,\alpha)$ as
@@
where $\mathbf{M}$ is the matrix of all pairwise log-ratios, of dimension $n\times\frac{m(m-1)}{2}$. Extensions to logistic regression and other generalized linear models are discussed in [6] and apply directly in our setting.
@@ 4.2 Pairwise log-ratio screening and separability index
    • For multiclass classification ($C>2$), we employ the generalized AUC proposed by [9],
@@ 4.3 Uncertainty quantification
  2. 2.
    Binary case ($C=2$). Two common approaches can be used to estimate $\operatorname{Var}(\widehat{AUC}_{jj^{\prime}})$.
    First, the classical approximation of [8]:
    $\operatorname{Var}_{\text{Hanley}}\!\bigl(\widehat{\mathrm{AUC}}_{jj^{\prime}}\bigr)$ $=\frac{AUC_{jj^{\prime}}(1-AUC_{jj^{\prime}})+(n_{j}-1)\bigl(Q_{1}-AUC_{jj^{\prime}}^{2}\bigr)+(n_{j^{\prime}}-1)\bigl(Q_{2}-AUC_{jj^{\prime}}^{2}\bigr)}{n_{1}n_{0}},$ (3)
    $Q_{1}$ $=\frac{AUC_{jj^{\prime}}}{2-AUC_{jj^{\prime}}},\;Q_{2}=\frac{2\,AUC_{jj^{\prime}}^{2}}{1+AUC_{jj^{\prime}}},$ (4)
    where $n_{j}$ and $n_{j^{\prime}}$ denote, respectively, the number of samples from class $j$ and class $j^{\prime}$ among the observations under comparison, that is, the samples being separated by the OTUs in $S$.
    Second, the nonparametric U-statistic approach of [5], for which a fast implementation is given by [12]. Define
@@
    where $AUC_{cc^{\prime}}$ is the binary AUC comparing classes $c$ and $c^{\prime}$. The variance of $AUC_{jj^{\prime}}$ is then obtained by propagating the uncertainty of its binary components:
  4. 4.
@@
    Here, $\operatorname{Var}(AUC_{cc^{\prime}})$ can be approximated using the binary variance formula in (3), whereas the covariances between binary pairs must be approximated or estimated; see below.
  5. 5.
    Dependence between AUCs across OTU pairs Each $AUC\_{jj^{\prime}}$ shares OTUs with other pairs, inducing dependence in the matrix $\mathbf{A}$. When aggregating these AUCs to compute the separability index $S_{k}$, this dependence propagates to its variance.
@@
    Pairs that share at least one OTU (e.g., $(AUC_{12},AUC_{13})$) are typically positively correlated. In practice, these covariances are non-negligible but difficult to derive analytically. A feasible approximation is The separability index $S_{k}$ is a linear combination of the pairwise AUCs within the set $S_{k}$, and its variance depends not only on the variances of individual AUCs, but also on their covariances. Exact expressions would require estimating a large number of covariance terms, which is computationally burdensome and unstable.
    To address this, we adopt a parsimonious approximation based on a single correlation parameter $\rho_{\text{OTU}}$ that captures the average dependence between AUCs sharing at least one common OTU. More precisely, we write
    $\text{Cov}\bigl(\widehat{\text{AUC}}_{jj^{\prime}},\widehat{\text{AUC}}_{kk^{\prime}}\bigr)\approx\begin{cases}\rho_{\text{OTU}}\,\sqrt{\widehat{\text{Var}}(\widehat{\text{AUC}}_{jj^{\prime}})\;\widehat{\text{Var}}(\widehat{\text{AUC}}_{kk^{\prime}})},&\text{if }\{j,j^{\prime}\}\cap\{k,k^{\prime}\}\neq\emptyset,\\ 0,&\text{otherwise}.\end{cases}$
    Substituting this approximation into the expression for $\text{Var}(S_{k})$ yields
    $\widehat{\text{Var}}(S_{k})=\mathbf{w}_{k}^{\top}\widehat{\Sigma}_{k}\mathbf{w}_{k},$
    where $\mathbf{w}_{k}$ is the vector of weights (all equal to 1/[k(k1)]1/[k(k-1)] in our case) and $\widehat{\Sigma}_{k}$ is the approximate covariance matrix of the AUCs associated with pairs in $S_{k}$.
    In practice, $\rho_{\text{OTU}}$ can be:
    • Fixed to a small positive value (e.g. 0.1–0.3) based on prior experience, or
    • Estimated from the data via a resampling scheme that compares empirical covariances of AUCs to the product of their standard errors.
    In our implementation we provide both options and allow the user to perform sensitivity analyses with respect to the choice of $\rho_{\text{OTU}}$. For the multiclass case, an additional layer of dependence arises because the Hand–Till AUC is itself an average of dependent binary pairs. An analogous approximation can be used for $\operatorname{Cov}(AUC_{cc^{\prime}},AUC_{ll^{\prime}})$ when the class pairs share at least one class.
    4.3.1 Alternative: non-parametric bootstrap.
    As an alternative to analytic variance formulas, we consider a non-parametric bootstrap procedure. This approach is particularly useful in capturing the sampling variability of $S_{k}$ under complex dependency structures:
@@
Overview of the Proposed Classification Framework for Microbiome Data Objective: Identify differential OTUs that best discriminate between groups $(g=1,\ldots,k)$ using compositional data analysis and pairwise log-ratio modeling. Input data: Abundance matrix $\mathbf{X}_{n\times m}$ where rows = samples, columns = OTUs, and class labels $Y\in\{1,\ldots,k\}$. 1. Data Preprocessing: